OpenStudy (anonymous):

Two negative integers have a sum of -12 and a product of 11. What are the integers?

OpenStudy (anonymous):

ab = 11 a + b = -12 use substitutionn: b = -12 - a a(-12 - a) = 11 so -a(12 + a) = 11 so 12a + a^2 = -11 so a^2 + 12a + 11 = 0 so (a + 11)(a + 1) = 0 so a = -11, -1 therefore b = -1, -11 thus the two numbers are -1 and -11

OpenStudy (anonymous):

Thanks.